OpenStudy (anonymous):

Find an equation of the tangent line to the graph of at the given point. MEDAL WILL BE REWARDED!!!! y=cos3x (pi/4, sqrt2/2)

OpenStudy (isaiah.feynman):

Find the derivative of y=cos3x, then substitute pi/4 for x, what ever you get is the SLOPE. Then using y-y1=m(x-x1) get the equation of the line by making the necessary substitutions!

OpenStudy (anonymous):

okay so is the derivative 0?

OpenStudy (anonymous):

or is it -3sin3x?

OpenStudy (isaiah.feynman):

Its -3sin3x.

OpenStudy (anonymous):

okay so my final answer is y=1, y=.9865?

OpenStudy (isaiah.feynman):

No y' = 0.123. That's the slope!

OpenStudy (anonymous):

so how do I plus that into the point slope form?

OpenStudy (isaiah.feynman):

And the slope is negative because of that minus sign.

OpenStudy (isaiah.feynman):

To put it in point slope form do the following substitution y1=sqrt2/2 m=-0.123, and x2=pi/4.